OpenStudy (anonymous):

5y3 – 4x3y2 = 6 implicitly derive

myininaya (myininaya):

derive means like to find so what do you want to find?

myininaya (myininaya):

i think you mean implicitly differentiate

OpenStudy (anonymous):

yes lol

myininaya (myininaya):

lol

myininaya (myininaya):

so what is (5y^3)'=?

OpenStudy (anonymous):

alright how to figure this out

OpenStudy (anonymous):

15y^2

myininaya (myininaya):

15y^2*y'

OpenStudy (amistre64):

remember to keep your derived bit around, they are needed later :)

myininaya (myininaya):

\[(4x^3y^2)'=4(x^3y^2)'=4(3x^2y^2+x^3(y^2)')\]

OpenStudy (amistre64):

they really should teach you to keep them at the start, and let you do the mental math in eliminating them if need be...

myininaya (myininaya):

what rule did i use there?

OpenStudy (anonymous):

chain rule

OpenStudy (amistre64):

...id say it begins with a pro, and ends with a duck, rule

myininaya (myininaya):

no i havent used the chain rule yet (but for (5y^3)' we did) we are looking at (4x^3y^2)' now

OpenStudy (anonymous):

doing online school for undergrad amistre dont trying to understand how to learn through these study groups before i head out again

OpenStudy (anonymous):

trying*

myininaya (myininaya):

right amistre product rule!

myininaya (myininaya):

but what is \[(y^2)'=?\]

myininaya (myininaya):

what rule do we need to apply here?

OpenStudy (anonymous):

haha guys are quick

OpenStudy (anonymous):

we need to apply chain rule

myininaya (myininaya):

yes so we do derivative of inside * derivative of outside y' * 2y^1 =y'2y

OpenStudy (anonymous):

5y3 – 4x3y2 = 6

OpenStudy (anonymous):

so derivative of 4x^2

myininaya (myininaya):

so anyways when we put it altogether we have \[15y^2y'-4(3x^2y^2+x^3y'2y)=0\]

OpenStudy (anonymous):

wolfram simplified it to 15y^2-8x^3y=0

myininaya (myininaya):

\[15y^2y'-12x^2y^2-8x^3yy'=0\] \[y'(15y^2-8x^3y)=12x^2y^2\] \[y'=\frac{12x^2y^2}{15y^2-8x^3y}\]